OpenStudy (anonymous):

Will give medal, please help!!!! How do I solve (k-7)^2 +6(k-7)+10 using the quadratic formula?

OpenStudy (fortytherapper):

I would first distribute and then combine like terms After that, we can use the quadratic formula

OpenStudy (anonymous):

okay but the answer is supposed to be 4 plus or minus imaginary i

OpenStudy (anonymous):

I tried using the substitution method but got a completely different answer

Parth (parthkohli):

hmm, I'm getting that. so you substituted \(x = k -7\) and got a quadratic in \(x\).

OpenStudy (anonymous):

yes

Parth (parthkohli):

\[x^2 + 6x + 10\]did you use the quadratic formula for this?

OpenStudy (anonymous):

when I substituted I got the x^2 +6x +10=0

OpenStudy (anonymous):

Then used that equation in the quadratic formula

Parth (parthkohli):

You should get \(x = -3 \pm i\). But remember that is not \(k\).

OpenStudy (anonymous):

But what do you do with that -3 plus or minus imaginary i then?

Parth (parthkohli):

You use \(x = k-7\) again.

Parth (parthkohli):

\[\Rightarrow k-7 = -3 \pm i\]

OpenStudy (anonymous):

Okay I got it thankyou

Parth (parthkohli):

No problem. :)